OpenStudy (anonymous):

Two boys who live 14 miles apart start at noon to walk toward each other at rates of 3 mph and 4 mph respectively. In how many hours will they meet?

OpenStudy (anonymous):

4 hours

OpenStudy (anonymous):

its actually 2 hours

OpenStudy (anonymous):

no its 4

OpenStudy (anonymous):

3x+4x=14 7x=14 x=14/7 x=2 hours they will meet.

OpenStudy (anonymous):

3+4=7 7/2 =3.5 14/3.5=4

OpenStudy (anonymous):

so who is right??

OpenStudy (anonymous):

click on the link

OpenStudy (anonymous):

i am

OpenStudy (anonymous):

but kaiden why would u do 7/2

OpenStudy (anonymous):

you take and add 3 and 4 then devide that by 2 to get the avegage speed

OpenStudy (anonymous):

u have to have a variable by 3 and 4 bc u dont know how many mph u go by so it has to be 3x+4x and total mph is 14

OpenStudy (anonymous):

i-need-help54 listen to me

OpenStudy (anonymous):

he is wrong